Amy is 28 years old and plans to start maximizing her 401(k) contributions offered through her employer - $19,000, starting this year.  

a.     Ignoring inflation and any company match, and assuming she is starting with a zero balance, how much will she have at age 60 assuming a 6% annual rate of return?

b.     Given the answer in part a and assuming Amy lives to 100, how much is the maximum amount she can spend on an annual basis starting at age 60. Assume she can earn a 6% annual rate of return during her retirement and inflation averages 3%.  

c.     Amy’s goal is to be able to withdraw $10,000 per month in retirement and she falls short if she just saves in her 401k account. What is another source of retirement income you may suggest that can help Amy reach her goal? Assume that there is no company match.  
Answer #1

Answer a)

Annual contribution (Pmt)= $ 19,000

Number of payment (nper) = starting from 28 years till 60 years = 33

Annual interest rate(r) = 6%

FV = FV ( r%, nper, pmt,)= FV(6%,33,-19000)= $ 184,9520.13

Answer b)

Net rate after retirement(r) = 6 -3=3%

Number of payment (nper) =starting from 60 years till 100 years =41

Accumulated corpus (PV) = $ 184,9520.13

Annual Payment (pmt) = pmt( rate, nper, pv,fv, type) =PMT(3%,41,--1849520.13,1)= $ 79,696.56

Answer c)  Amy can use either all or any of them from suggested below list to reach her retirement goal.

  • Roth IRAs
  • Simplified Employee Pension Individual Retirement Account (SEP IRA's)
  • A variable annuity :(contract for investor and an insurance company.)
  • Index funds.